OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write a system of two equations with the variables m and f in each to represent this problem. pls dont goggle, answer it from what you know :) thx

OpenStudy (hoblos):

40= m + 10f 25= m + 5f 15 = 5f ---> f=3 25= m+15-----> m=10